UT 20170925 LUCI1/2 AO Nighttime Commissioning

Observer LUCI-Team: JHeidt, DThompson (LBTO)
Observer AO-Team: DMiller, GTaylor (LBTO)
Telescope Operator: Sallenson

Observations done remotely from Tucson

Plan for the night

- LUCI-AO short images Bin=1

- LUCI-AO AFC @ DX

Summary

LUCI-AO imaging Bin = 1 / 400 modes works

First successfull AFC-tests JHKs @ low elevation

Failure rate of the AO-system(s) is about 1 / hour

Details

All times are UT
Files are luci1/2.20170925.NNNN.fits

01:20 Sunst, clear skies, T = +5 deg, hum = 40%, wind ~ 5 m/s

ALTA predicts good seeing all night

01:25 Steve opens up, checks pointing and collimation

Bin = 1, binocular mode, 400 modes reconstructor

01:51 Preset to P Cygni, Fe II and Br_gam, OK, RunAO, OK, loops closed

02:02 Error on offset @ DX (0",0"), arbitrator died, ...what a start into the night...
restart Adec @ DX

02:23 Resent preset, OK, RunAO, OK

02:41 Starting with large offsets, when coming back to target, target wasn't back on the AO-pyramid
recovered manually

DIMM 1" - 1.2"

03:33 Loop opened @ SX after large offset (30" within the Patrol field) due to a Pupil shift at the AO-system)
this is a known issue (IT #6333)

DIMM 1" - 1.8" th high values occured while passing meridian at 5deg from zenith, ie
its rather an issue of the DIMM itself

04:09 Offset failed @ SX, reason unknown (perhaps scope moved before loop was open?)

04:19 Observations resumed,

04:53 Loop opened @ DX, reason unknown, recovered manually, cont'd

L1: # 2 - 101
L2: # 2 - 101

That wasn't that bad !

AFC-ima @ DX

Steve authorzies LUCI2 only

05:40 Preset to M15, OK, RunAO, OK,

05:53 GUIs hang after acquisition image of M15, XKILL + restart OBSGUI

Acquisition #102-104

Ks: AFC #105 (short image), 50 pixels off, how can that be?
# 106 300sec image, off the holes
# 107,108 redo alignment
# 109, short image OK
# 110 - 112 SRR, images don't look good
# 113-117 short images

06:10 XKILL required again

06:22 XKILL once more, logged in at obs2 at the mountain per remote desktop and continue from there

looking into the (low) wind causing vibrations, thats why the images are not so good

06:40 Moving to Palomar 1

Steve checks pointing before as it is a long slew

06:45 Preset to star 1 in Palomar 1, OK, RunAO, processes stuck, failed , resent preset, RunAO failed again
(Wunit became disconnected for whatever reason)

07:04 Preset to Palomar 1 again, failed, AOS lost elevation and wind , shell had to be reset,

07:15 Preset to Palomar 1 once more, OK, RunAO failed (too much automatic gain)

07:24 Preset to Palomar 1 again, OK, RunAO OK,

Bin = 3

K: acquisition, # 118 - 120

K: SRR 10sec, way off as in M15

Ks: 300 sec, #121 dito

redo

K: # 123-126

K: # 127, SRR 10sec, way off again ! Why ???

Ks: 300 sec, #128 dito

Checked that with Dave and it worked, # 129-132

Check this again, resent preset,

Dave know the answer: preset is done without passive flexure compensation, but must be !

08:17 Modified script, resent preset to verify, # 1331 - 140

08:28 Preset to Mayall 2, failed, Steve corrects pointing, resent preset, OK, RunAO,

Bin = 4

DIMM = around 1"

K: Acquisition, # 141 - 142

Ks: # 143 - 145 (300sec, SRR), 146-150 (1min, LIR)

About 6pix FWHM on SRR-images

K: Acquisition, #152 - 154

H: # 155 - 157 (300sec, SRR), 1st 2 images OK, 3rd image AFC stopped working

H: # 158 - 162 (1min, LIR)

Redo H-band:

K: Acquisition # 163 - 165

H: # 166 same issue as in ima # 157, corrections in x are getting larger and larger until they are too large

Problem seems to be new positions of the spots as the detector was new aligned in Summer with
a change of 7deg, Dave measured new positions and Alexander put them in. In addition
one channel doesn't work properly even more impacting the AFC.

Taking some Ks-images in the meantime 6*10sec integrated, LIR, offsets done manually, # 162 - 176

11:23 Resume AFC-testing

K: Acquisition, # 183 - 185

H: SRR, 300sec, # 186 - 188, worked this time smile

J: SRR, 300sec, # 189 - 191, worked also, # 192 - 196

DIMM around 1" during the last 3hrs

12:10 Handed over to Dave for ESM-tests

End of run

-- %USERSIG{JochenHeidt - 2017-09-25}%
Topic revision: r6 - 25 Sep 2017, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback